Can You Please Solve This Question:
find three consecutive terms of an a.p whoose sum is 30 and product is 960

Answer by Alan3354(69443)

Can You Please Solve This Question:
find three consecutive terms of an a.p whoose sum is 30 and product is 960
-----------
(x-n) + x + (x+n) = 30
3x = 30
x = 10
-----------
(x-n)*x*(x+n) = 960
(10-n)*10*(10+n) = 960
(10-n)*(10+n) = 96
100 - n^2 = 96
n = 2
--> 8, 10 & 12